OpenStudy (anonymous):

solve: 3(5-y)=3(y+4)

OpenStudy (anonymous):

3(5-y)/3=3(y+4)/3

OpenStudy (anonymous):

5-y=y+4

OpenStudy (anonymous):

5-4=2y 1/2=2y/2

OpenStudy (anonymous):

y=1/2